OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(2) + g(3). is the answer 4 this problem (15)?????

OpenStudy (anonymous):

I read as follows \[f(x) = x^2 + 1 \] and \[g(x) = 3x + 1\] then \[f(2) = (2^2) + 1\] and \[g(3) = 3*3 + 1\] solve both equations and find f(2) + g(3) to get you answer

OpenStudy (anonymous):

ok i tried f=5 and g=10 what am i suppose to do next just 5+10 and that =15...is that right i have done this worksheet 3 times...